OpenStudy (anonymous):

y varies jointly with p and q. When p = 5 and q = 2, y = 80. Find y when p = 4 and q = 5.

jimthompson5910 (jim_thompson5910):

"y varies jointly with p and q" means that y = kpq for some (not known yet) constant k

jimthompson5910 (jim_thompson5910):

"When p = 5 and q = 2, y = 80" this means that y = kpq becomes 80 = k*5*2 which turns into 80 = 10k

jimthompson5910 (jim_thompson5910):

Solve that for k and tell me what you get

OpenStudy (anonymous):

8?

jimthompson5910 (jim_thompson5910):

yes good

jimthompson5910 (jim_thompson5910):

So y = kpq turns into y = 8pq

OpenStudy (anonymous):

so its 8*4*5

jimthompson5910 (jim_thompson5910):

Last part: Find y when p = 4 and q = 5 Plug in p = 4 and q = 5 to go from y = 8pq to y = 8*4*5 y = ??

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcoem

jimthompson5910 (jim_thompson5910):

welcome*